# Efficient calculation of magnetic fields from ferromagnetic materials near strong electromagnets, and application to stellarator coil optimization

### v1

@@ 1 Introduction
In the blanket of a fusion reactor, the leading candidates for structural materials are steels that are ferromagnetic [tanigawa2017development, mergia2008structural, hirose2014physical]. This choice is based on the requirements that the material maintain adequate structural properties and low nuclear activation in the presence of high neutron fluxes and high temperatures. Ferromagnetic materials are also of interest in fusion for reducing magnetic field ripple from toroidal field coils [turner1978, sheffield1993use, kawashima2001demonstration]. In fusion and in other magnetic systems such as magnetic resonance imaging (MRI), pieces of ferritic material can be used as shims to fine-tune the field [dorri2002passive, lopez2008passive]. As physics properties of the confined plasma can be highly sensitive to the magnetic field, it is critical to compute the magnetic field including the effects of the magnetic materials, which have motivated studies for both tokamaks [tsuzuki2003high, tobita2003reduction, shinohara2011effects, chiariello2015effective] and stellarators [ikuta1983helical, harmeyer1999effect, ji2017investigation]. For all these applications, it is desirable to have a fast and flexible method to compute the magnetic field including the ferromagnetic materials, to include in physics calculation workflows. Moreover, for design optimization, where the field must be evaluated hundreds of times or more, efficiency is a priority, and it is valuable to have derivatives of the field with respect to design variables. In the present work, we propose a new calculation method that satisfies these criteria, and demonstrate its use for shape and current optimization of stellarator electromagnets.
@@
A fundamental challenge in calculating the field in the presence of ferromagnetic materials is that magnetic material in one region affects the magnetization in every other region. Therefore to achieve consistency, it is necessary to solve a large system of equations. Furthermore, the relationship between field and magnetization is generally nonlinear, making this system of equations nonlinear, requiring iteration to solve. Approaches include solving for the field [newman1972gfun], scalar magnetic potential [carpentier2013resolution], or vector potential [le2015magnetic] in each element, or using analytic formulas for the field of polyhedra used for the mesh [bjork2021magtense]. These methods share the advantage that there is no need to mesh the volume outside the magnetic material; only the ferromagnetic domain is meshed.
@@
In contrast to these methods, our approach eliminates the need to solve any system of equations and does not require any iteration. This simplification is made possible by making two approximations which are well satisfied for fusion systems, as we will show. The first approximation is that the ferromagnetic materials perturb the total field by a small amount compared to the field from the electromagnets and plasma current. In several examples shown in the following sections, this approximation is verified to be accurate. Second, as the saturation magnetization of the material is achieved at field strengths much smaller than the typical fields in magnetic fusion, the material is saturated. This is true especially for high-field reactor designs based on high-temperature superconductors [creely2020overview, hegna2025infinity, lion2025stellaris], but is true also for lower-field designs based on low-temperature superconductors. Together, these approximations mean that the magnetization vector has a known constant magnitude, pointing along the local field computed in the absence of the steel. Since in this model ferromagnetic material in one region does not affect magnetization in another, no system of equations needs to be solved, and the field can be computed rapidly by straightforward evaluation of Biot-Savart-type integrals without meshing the volume outside the magnetic material. Specifically our method uses the field from a large number of point dipoles, which conveniently has already been available in several stellarator design frameworks including SIMSOPT [landreman2021simsopt] and FAMUS [zhu2020topology], so minimal software development was required to implement the method.
@@
Our method differs from previous work on magnetic materials in fusion in several ways. Our approach is most similar to the one in the code FEMAG [urata2003, shinohara2011effects], which also makes the same two approximations described above. However the FEMAG approach also requires that the magnetic material is represented as infinitesmally thin and flat rectangular plates, whereas the approach here allows for arbitrary computer-aided design (CAD) geometry. Our approach also differs from the previous stellarator analysis in [harmeyer1999effect] which assumed a spatially uniform permeability; that is not an accurate approximation since in fact the magnetization has uniform magnitude due to saturation whereas the field magnitude varies considerably, so their ratio is nonuniform. Our work can also be contrasted with the related problem of producing stellarator fields with hard permanent magnets [helander2020stellarators, zhu2020topology, hammond2022design, kaptanoglu2023greedy]. The permanent magnet problem considered in those references is fundamentally different as the magnetization was considered to be insensitive to the field from electromagnetic coils and plasma current, in contrast to the case of soft ferromagnetic materials considered here.
@@
Our mathematical method is described in detail in the following section. To verify its accuracy, a benchmark with the commercial finite-element code COMSOL [comsol] is presented in section 3, where excellent agreement is demonstrated for the geometry of a modern quasi-isodynamic stellarator reactor. The coil optimization example is then presented in section 4. In that section, we also demonstrate how the effects of materials can be included straightforwardly in free-boundary magnetohydrodynamic (MHD) equilibrium calculations, to assess the impact on physics properties such as confinement and stability. Overall we find that the magnetic materials have little effect on the plasma physics properties of the self-consistent MHD equilibria, with the largest effects being a reduction in rotational transform [harmeyer1999effect, ji2017investigation] that causes a shift in the edge magnetic islands, and a slight destabilization of ideal ballooning modes. Both effects are compensated by the coil reoptimization that includes the magnetic materials. Finally in section 5 we discuss the results and conclude.
@@ 2.1 Approximations and model equations
To evaluate the magnetization, consider the MHM-H curve for a typical steel under consideration for fusion blankets, EUROFER97 [mergia2008structural], shown in figure 1. Converting the $H$ values on the x-axis to Tesla by multiplying by $\mu_{0}$, the range of the x-axis corresponds to $\pm 1.26$ Tesla. It is apparent that the magnetization saturates by $\lesssim 0.3T$. Typical field strengths in the blanket region of fusion reactor designs are far higher than this, several Tesla or more. Thus, the magnitude $M=|\mathbf{M}|$ will be fixed at the saturation value, which we denote $M_{sat}$. The magnetization is not linearly related to the field so one should not make an approximation of linear permeability.
Figure 1: MHM-H curve for EUROFER97. Figure reproduced from [mergia2008structural]. The x-axis range of $\pm 1000$ kA/m is equivalent to $\pm 1.26$ T.
The saturation magnetization of EUROFER97 is $M_{sat}\sim 1.4\times 10^{6}$ A/m, so $\mu_{0}M_{sat}\sim 1.7$ T. The value for F82M steel is nearly identical [hirose2014physical]. This value again is small compared to typical reactor field strengths, particularly for recent high-field designs, so we can approximate $\mu_{0}|\mathbf{M}|\ll|\mathbf{B}|$ and $|\mathbf{M}|\ll|\mathbf{H}|$.
@@
The quantity $\mathbf{B}^{plasma}$ can be computed from the Biot-Savart formula by integrating the plasma current over the plasma volume, or by an integral over the plasma boundary via the virtual casing principle [shafranov1972use, drevlak2005pies]. Note that the evaluation point for $\mathbf{B}^{plasma}$ is strictly outside the plasma, so there is no singular denominator, and sophisticated methods for virtual casing evaluation on the plasma boundary [malhotra2019efficient] are not required. In sum, the magnitude and direction of $\mathbf{M}$ are known, so eq (1) can be evaluated.
@@ 2.2 Quadrature
We must integrate eq (1) over each volume containing magnetic material. For numerical calculations, this requires a set of discrete quadrature points $\mathbf{r}_{k}$ and weights $w_{k}$, such that volume integrals become $\int_{\Omega}d^{3}r\to\sum_{k}\mathbf{r}_{k}w_{k}$. The quadrature points effectively represent point dipoles. Here we describe two methods for generating these quadrature grids, using unstructured or structured grids.
The unstructured grid approach is convenient for interfacing with CAD designs, and is required for components with holes such as ports. Given a CAD description of the steel component, a 3D unstructured mesh is generated. The barycenter of each mesh element then gives $\mathbf{r}_{k}$, and the element volume gives the weight $w_{k}$.
@@
In either the unstructured or structured approach, for calculations to be accurate, the distance between quadrature points within the ferromagnetic region should be smaller than the distance to the evaluation point. The latter is typically the distance to the plasma. Thus, for components that are closest to the plasma, finer mesh resolution is required. If the field must be computed very close to the magnetic material, the formula for the field from a uniformly magnetized tetrahedron [nielsen2019stray] can be applied, with the magnetization computed at the element barycenter as in the dipole method. However we have found that evaluating the field from the simpler dipole formula is sufficient for the applications here.
@@ 2.3 Forces
The method here can also be used to rapidly compute the magnetic force on ferromagnetic components. The force on a point dipole $\mathbf{m}$ is $\mathbf{f}=(\nabla\mathbf{B}^{ext})\cdot\mathbf{m}$ [griffiths2023introduction], where $\mathbf{B}^{ext}$ is the field due to sources other than the dipole, so the force density on a magnetized region is $\mathbf{F}=(\nabla\mathbf{B}^{ext})\cdot\mathbf{M}$. This same formula can also be derived by summing the Lorentz force densities for the volume and boundary of the ferromagnetic region, $\mathbf{F}=\int_{\Omega}d^{3}r\mathbf{J}_{b}\times\mathbf{B}^{ext}+\int_{\partial\Omega}d^{2}r\mathbf{K}_{b}\times\mathbf{B}^{ext}$, where $\mathbf{J}_{b}=\nabla\times\mathbf{\mathbf{M}}$ is the current density associated with bound currents and $\mathbf{K}_{b}=\mathbf{M}\times\mathbf{n}$ is the surface current density associated with bound currents. This equivalence and other approaches to computing the force are discussed further in the appendix. The total force on a component follows as
@@ 3 Verification with the finite element method
To verify that our approximate model is accurate for fusion reactor parameters, we carry out a comparison with the commercial finite-element software COMSOL Multiphysics®[comsol]. The geometry for the comparison is shown in figure 3. The plasma shape, shown in pink, is a three-field-period quasi-isodynamic stellarator. The configuration is at reactor scale, with average major radius of 12.7 m, and volume-averaged field strength in the plasma of 8 T. The plasma is confined by a set of 36 modular coils (6 unique shapes each repeated 6 times), shown in gray. While multi-filament finite-build coils are being used for the design of Type One Energy’s stellarators, a single-filament model of the coils is used in this example for simplicity; use of multi-filament coils does not alter any aspect of the method. Two walls of EUROFER97 are placed between the plasma and coils, one 60 cm from the plasma, and the other 140 cm from the plasma, both 15 cm thick, with holes removed for ports. These components are not from an actual blanket design, but are merely intended to illustrate a possible application of the method. Only one half period of the walls is shown in the right two panels, but they extend around the full torus, as in the left panel. Considering only the field from the coils, the minimum field strength in the plasma region is >6.4>6.4 T, and the minimum in the steel volume is >4.3>4.3 T, so the steel is fully saturated. However the COMSOL calculation uses the full $M(H)$ curve for EUROFER97 without assuming saturation. COMSOL also fully accounts for the effect of magnetization in one region on the magnetization in another. Calculations with the approximate model are performed by first meshing the steel CAD geometry using GMSH [geuzaine2009gmsh], then applying the open-source SIMSOPT software [landreman2021simsopt], using its DipoleField class for arrays of point dipoles. Plasma current is neglected for simplicity in this COMSOL comparison, but will be considered in later sections.
@@
The total field including the steel is computed using both the approximate model and COMSOL. The field on the plasma boundary is displayed in figure 4. The four panels in the top-left show the field from only the coils, $\mathbf{B}^{coils}$, and the total field $\mathbf{B}^{total}=\mathbf{B}^{coils}+\mathbf{B}^{ferr}$, for both the model and COMSOL. Note that the color scale for these four panels is identical. These four panels are visually indistinguishable, showing both that the effect of the steel is small, and that there is good agreement between the two methods. The remaining four panels show differences between these fields on a much finer color scale so they can be seen, using the same color scale for the four panels. The bottom row shows the difference between our model and COMSOL both for the coil field alone and for the total field. The differences in coil field alone (bottom left) are due only to discretization error since the underlying equations are identical in the absence of steel. The differences in total field (bottom center) arise due to both discretization error and the different underlying equations. However the discrepancy in the bottom center panel is not significantly larger than the bottom left panel, indicating that the physics differences are not larger than discretization error. The right two panels show $|\mathbf{B}^{ferr}|$ computed by the two approaches, showing close similarity. The maximum $|\mathbf{B}^{ferr}|$ occurs under the vertical edges of the ports.
@@
We can also directly assess the accuracy of the approximation used to determine the magnetization in the reduced model. This is done in figure 5, which shows the norm of the difference in magnetization vectors between COMSOL and the reduced model. The magnetization is evaluated on a surface 0.675 m from the plasma, in the middle of the inner steel layer. The difference is normalized to the magnitude of the magnetization in the reduced model, i.e. the saturation magnetization. It can be seen that this relative difference in magnetization is no more than a few percent, meaning the approximation is quite good.
@@
The forces on the steel can also be compared between the two methods. In COMSOL the force is calculated from a surface integral of the Maxwell stress on the boundary instead of the volume integral (6). For this comparison we evaluate the net force on one half field period of the inner steel region. Results are displayed in figure 6. The Cartesian components of this force in MN were found to be [-22.1, -24.2, -22.3] from COMSOL, and [-21.8, -24.0, -21.9] according to the reduced method introduced in this article, based on eq 6). Thus, good agreement is obtained.

### v2

@@ 1 Introduction
In the blanket of a fusion reactor, the leading candidates for structural materials are steels that are ferromagnetic [42, 34, 20]. This choice is based on the requirements that the material maintain adequate structural properties and low nuclear activation in the presence of high neutron fluxes and high temperatures. Ferromagnetic materials are also of interest in fusion for reducing magnetic field ripple from toroidal field coils [45, 38, 26]. In fusion and in other magnetic systems such as magnetic resonance imaging (MRI), pieces of ferritic material can be used as shims to fine-tune the field [11, 31]. As physics properties of the confined plasma can be highly sensitive to the magnetic field, it is critical to compute the magnetic field including the effects of the magnetic materials, which have motivated studies for both tokamaks [44, 43, 39, 7] and stellarators [22, 17, 24]. For all these applications, it is desirable to have a fast and flexible method to compute the magnetic field including the ferromagnetic materials, to include in physics calculation workflows. Moreover, for design optimization, where the field must be evaluated hundreds of times or more, efficiency is a priority, and it is valuable to have derivatives of the field with respect to design variables. In the present work, we propose a new calculation method that satisfies these criteria, and demonstrate its use for shape optimization of stellarator electromagnets.
@@
A fundamental challenge in calculating the field in the presence of ferromagnetic materials is that magnetic material in one region affects the magnetization in every other region. Therefore to achieve consistency, it is necessary to solve a large system of equations. Furthermore, the relationship between field and magnetization is generally nonlinear, making this system of equations nonlinear, requiring iteration to solve. Approaches include solving for the field [35], scalar magnetic potential [6], or vector potential [29] in each element, or using analytic formulas for the field of polyhedra used for the mesh [3]. These methods share the advantage that there is no need to mesh the volume outside the magnetic material; only the ferromagnetic domain is meshed.
@@
In contrast to these methods, our approach eliminates the need to solve any system of equations and does not require any iteration. This simplification is made possible by making two approximations which are well satisfied for fusion systems, as we will show. The first approximation is that the ferromagnetic materials perturb the total field by a small amount compared to the field from the electromagnets and plasma current. In several examples shown in the following sections, this approximation is verified to be accurate. Second, as the saturation magnetization of the material is achieved at field strengths much smaller than the typical fields in magnetic fusion, the material is saturated. This is true especially for high-field reactor designs based on high-temperature superconductors [9, 18, 30], but is true also for lower-field designs based on low-temperature superconductors. Together, these approximations mean that the magnetization vector has a known constant magnitude, pointing along the local field computed in the absence of the steel. Since in this model ferromagnetic material in one region does not affect magnetization in another, no system of equations needs to be solved, and the field can be computed rapidly by straightforward evaluation of Biot-Savart-type integrals without meshing the volume outside the magnetic material. Specifically our method uses the field from a large number of point dipoles, which conveniently has already been available in several stellarator design frameworks including SIMSOPT [28] and FAMUS [47], so minimal software development was required to implement the method.
@@
Our method differs from previous work on magnetic materials in fusion in several ways. Our approach is most similar to the one in the code FEMAG [46, 39], which also makes the same two approximations described above. However the FEMAG approach also requires that the magnetic material is represented as infinitesmally thin and flat rectangular plates, whereas the approach here allows for arbitrary computer-aided design (CAD) geometry. Our approach also differs from the previous stellarator analysis in [17] which assumed a spatially uniform permeability; that is not an accurate approximation since in fact the magnetization has uniform magnitude due to saturation whereas the field magnitude varies considerably, so their ratio is nonuniform. Our work can also be contrasted with the related problem of producing stellarator fields with hard permanent magnets [19, 47, 16, 25]. The permanent magnet problem considered in those references is fundamentally different as the magnetization was considered to be insensitive to the field from electromagnetic coils and plasma current, in contrast to the case of soft ferromagnetic materials considered here.
@@
Our mathematical method is described in detail in the following section. To verify its accuracy, a benchmark with the commercial finite-element code COMSOL [8] is presented in section 3, where excellent agreement is demonstrated for the geometry of a modern quasi-isodynamic stellarator reactor. The coil optimization example is then presented in section 4. In that section, we also demonstrate how the effects of magnetic materials can be included straightforwardly in free-boundary magnetohydrodynamic (MHD) equilibrium calculations, to assess the impact on physics properties such as confinement and stability. Overall we find that the magnetic materials have little effect on the plasma physics properties of the self-consistent MHD equilibria, with the largest effects being a reduction in rotational transform [17, 24] that causes a shift in the edge magnetic islands, and a slight destabilization of ideal ballooning modes. Both effects are compensated by the coil reoptimization that includes the magnetic materials. Finally in section 5 we discuss the results and conclude.
@@ 2.1 Approximations and model equations
To evaluate the magnetization, consider the MHM-H curve for a typical steel under consideration for fusion blankets, EUROFER97 [34], shown in figure 1. Converting the $H$ values on the x-axis to Tesla by multiplying by $\mu_{0}$, the range of the x-axis corresponds to $\pm 1.26$ Tesla. It is apparent that the magnetization saturates by $\lesssim 0.3T$. Typical field strengths in the blanket region of fusion reactor designs are far higher than this, several Tesla or more. Thus, the magnitude $M=|\mathbf{M}|$ will be fixed at the saturation value, which we denote $M_{sat}$. The magnetization is not linearly related to the field so one should not make an approximation of linear permeability.
Figure 1: MHM-H curve for EUROFER97. Figure reproduced from [34]. The x-axis range of $\pm 1000$ kA/m is equivalent to $\pm 1.26$ T.
The saturation magnetization of EUROFER97 is $M_{sat}\sim 1.4\times 10^{6}$ A/m, so $\mu_{0}M_{sat}\sim 1.7$ T. The value for F82M steel is nearly identical [20]. This value again is small compared to typical reactor field strengths, particularly for recent high-field designs, so we can approximate $\mu_{0}|\mathbf{M}|\ll|\mathbf{B}|$ and $|\mathbf{M}|\ll|\mathbf{H}|$.
@@
The quantity $\mathbf{B}^{plasma}$ can be computed from the Biot-Savart formula by integrating the plasma current over the plasma volume, or by an integral over the plasma boundary via the virtual casing principle [37, 12]. Note that the evaluation point for $\mathbf{B}^{plasma}$ is strictly outside the plasma, so there is no singular denominator, and sophisticated methods for virtual casing evaluation on the plasma boundary [32] are not required. In sum, the magnitude and direction of $\mathbf{M}$ are known, so eq (1) can be evaluated.
@@ 2.2 Quadrature
We must integrate eq (1) over each volume containing magnetic material. For numerical calculations, this requires a set of discrete quadrature points $\mathbf{r}_{k}$ and weights $w_{k}$, such that volume integrals become $\int_{\Omega}d^{3}r\,f(\mathbf{r})\to\sum_{k}\,w_{k}f(\mathbf{r}_{k})$. The quadrature points effectively represent point dipoles. Here we describe two methods for generating these quadrature grids, using unstructured or structured grids.
The unstructured grid approach is convenient for interfacing with CAD designs, and is required for components with holes such as ports. Given a CAD description of the steel component, a 3D unstructured mesh is generated. The barycenter of each mesh element then gives $\mathbf{r}_{k}$, and the element volume gives the weight $w_{k}$. Discretization error for this integration rule scales as $\sim O(h^{2})$ for element size $h$ [48]. Higher order quadrature rules over common element types are known and could be used if desired.
@@
In either the unstructured or structured approach, for calculations to be accurate, the distance between quadrature points within the ferromagnetic region should be smaller than the distance to the evaluation point. The latter is typically the distance to the plasma. Thus, for components that are closest to the plasma, finer mesh resolution is required. If the field must be computed very close to the magnetic material, the formula for the field from a uniformly magnetized tetrahedron [36] can be applied, with the magnetization computed at the element barycenter as in the dipole method. However we have found that evaluating the field from the simpler dipole formula is sufficient for the applications here.
@@ 2.3 Forces
The method here can also be used to rapidly compute the magnetic force on ferromagnetic components. The force on a point dipole $\mathbf{m}$ is $\mathbf{f}=(\nabla\mathbf{B}^{ext})\cdot\mathbf{m}$ [15], where $\mathbf{B}^{ext}$ is the field due to sources other than the dipole, so the force density on a magnetized region is $\mathbf{F}=(\nabla\mathbf{B}^{ext})\cdot\mathbf{M}$. This same formula can also be derived by summing the Lorentz force densities for the volume and boundary of the ferromagnetic region, $\mathbf{F}=\int_{\Omega}d^{3}r\mathbf{J}_{b}\times\mathbf{B}^{ext}+\int_{\partial\Omega}d^{2}r\mathbf{K}_{b}\times\mathbf{B}^{ext}$, where $\mathbf{J}_{b}=\nabla\times\mathbf{\mathbf{M}}$ is the current density associated with bound currents and $\mathbf{K}_{b}=\mathbf{M}\times\mathbf{n}$ is the surface current density associated with bound currents. This equivalence and other approaches to computing the force are discussed further in the appendix. The total force on a component follows as
@@ 3 Verification with the finite element method
To verify that our approximate model is accurate for fusion reactor parameters, we carry out a comparison with the commercial finite-element software COMSOL Multiphysics®[8]. The geometry for the comparison is shown in figure 3. The plasma shape, shown in pink, is a three-field-period quasi-isodynamic stellarator. The configuration is at reactor scale, with average major radius of 12.7 m, and volume-averaged field strength in the plasma of 8 T. The plasma is confined by a set of 36 modular coils (6 unique shapes each repeated 6 times), shown in gray. While multi-filament finite-build coils are being used for the design of Type One Energy’s stellarators, a single-filament model of the coils is used in this example for simplicity; use of multi-filament coils does not alter any aspect of the method. Two walls of EUROFER97 are placed between the plasma and coils, one 60 cm from the plasma, and the other 140 cm from the plasma, both 15 cm thick, with holes removed for ports. These components are not from an actual blanket design, but are merely intended to illustrate a possible application of the method. Only one half period of the walls is shown in the right two panels, but they extend around the full torus, as in the left panel. Considering only the field from the coils, the minimum field strength in the plasma region is >6.4>6.4 T, and the minimum in the steel volume is >4.3>4.3 T, so the steel is fully saturated. However the COMSOL calculation uses the full $M(H)$ curve for EUROFER97 without assuming saturation. COMSOL also fully accounts for the effect of magnetization in one region on the magnetization in another. The default second-order elements are used. Calculations with the approximate model are performed by first meshing the steel CAD geometry using GMSH [14], then applying the open-source SIMSOPT software [28], using its DipoleField class for arrays of point dipoles. Plasma current is neglected for simplicity in this COMSOL comparison, but will be considered in later sections.
@@
The total field including the steel is computed using both the approximate model and COMSOL. The field on the plasma boundary is displayed in figure 4. The four panels in the top-left show the field from only the coils, $\mathbf{B}^{coils}$, and the total field $\mathbf{B}^{total}=\mathbf{B}^{coils}+\mathbf{B}^{ferr}$, for both the model and COMSOL. Note that the color scale for these four panels is identical. These four panels are visually indistinguishable, showing both that the effect of the steel is small, and that there is good agreement between the two methods. The remaining four panels show differences between these fields on much finer color scales so they can be seen. The right pair of panels share a common color scale, and the bottom pair of panels share a different common color scale. The bottom row shows the difference between our model and COMSOL both for the coil field alone and for the total field. The differences in coil field alone (bottom left) are due only to discretization error since the underlying equations are identical in the absence of steel. The differences in total field (bottom center) arise due to both discretization error and the different underlying equations. However the discrepancy in the bottom center panel is not much larger than the bottom left panel, indicating that the physics differences are comparable to discretization error. The right two panels show $|\mathbf{B}^{ferr}|$ computed by the two approaches, showing close similarity. The maximum $|\mathbf{B}^{ferr}|$ occurs under the vertical edges of the ports.
@@
We can also directly assess the accuracy of the approximation used to determine the magnetization in the reduced model. This is done in figure 5, which shows the norm of the difference in magnetization vectors between COMSOL and the reduced model. The magnetization is evaluated on a surface 0.675 m from the plasma, in the middle of the inner steel layer. In the ports, visible as the rectangular regions on the top and bottom left of the figure, the difference is necessarily zero. The difference is normalized to the magnitude of the magnetization in the reduced model, i.e. the saturation magnetization. It can be seen that this relative difference in magnetization is no more than a few percent, meaning the approximation is quite good.
@@
The forces on the steel can also be compared between the two methods. In COMSOL the force is calculated from a surface integral of the Maxwell stress on the boundary instead of the volume integral (6). For this comparison we evaluate the net force on one half field period of the inner steel region. Results are displayed in figure 6. The Cartesian components of this force in MN were found to be [-22.1, -24.2, -22.3] from COMSOL, and [-21.8, -24.0, -21.9] according to the reduced method introduced in this article, based on eq 6). Thus, good agreement is obtained. Agreement is also found for the volumetric force density within the steel, the magnitude of which is shown in figure 7. As with figure 5, data are shown in the middle of the inner steel layer. Of the various approaches that can be considered to compute magnetic forces (A), the Kelvin formula $\mu_{0}\mathbf{M}\cdot\nabla\mathbf{H}$ [33, 4] is used to compute the force density in COMSOL. This quantity is somewhat noisy since the finite element solution for the magnetic field must be numerically differentiated. However COMSOL and the approximate model generally show agreement in the magnitude and spatial pattern of the force density.
@@
Figure 7: The approximate model presented in this article also provides the force density within steel components, showing general agreement with COMSOL.